Please see question below. Please double check all work and units/unit conversions. Show all work. Thank you!

2a) A charge of -2.8 micro-coulombs lies at the origin and another charge of +9.5 micro- coulombs lies at x--14 cm, y +4 cm. At a certain point, the total electric field from these two charges is zero. What is the y coordinate of this point in cm? Include a negative sign if negative. 2b) A charge of +8 micro-coulombs is placed at x-0 cm, y +17 cm, a charge of -5 micro- coulombs is placed at x-0 cm, y -24 cm, a charge of +9 micro-coulombs is placed at x -7 cm, y 0 cm, and an unknown charge electric field at the origin is 298 degrees counter-clockwise from the positive x axis. What is the unknown charge in micro-coulombs? If negative, include a negative sign in your answer is placed! at x +25 cm, y = 0 cm. The angle of the total
Answer #1

2a)

Q2 = 9.5 x 10-6 C 0.14 m 2.8 x 106 C I d Sin E1 E2

consider the triangle AOB

using Pythagorean theorem

r = distance between the two charges = AB = sqrt(OA2 + OB2) = sqrt(0.042 + 0.142) = 0.15 m

heta = tan-1(OA/OB) = tan-1(0.04/0.14) = 15.94 deg

consider point C

for the electric field at C to be zero

E1 = E2

k q1/d2 = k q2/(r + d)2

q1/d2 = q2/(r + d)2

(2.8 x 10-6)/d2 = (9.5 x 10-6)/(0.15 + d)2

d = 0.18 m

y-coordinate : - d Sinheta = - (0.18) Sin15.94 = - 0.0494 m = - 4.9 cm

2 b)

g乂! 6 2 017m

when q4 is positive

Net electric field at the origin along X-direction is given as

Ex = E3 - E4 = k q3/r32 - k q4/r42 = (9 x 109) (9 x 10-6)/(0.07)2 - (9 x 109) q4/(0.25)2

Net electric field at the origin along Y-direction is given as

Ey = E1 + E2 = k q1/r12 + k q2/r22 = (9 x 109) (8 x 10-6)/(0.17)2 + (9 x 109) (5 x 10-6)/(0.24)2 = 3.3 x 106

Given that :

tanheta = Ey /Ex

tan62 = ( 3.3 x 106)/Ex

Ex = 1.8 x 106 N/C

when q4 is positive

Ex = E3 - E4 = k q3/r32 - k q4/r42 = (9 x 109) (9 x 10-6)/(0.07)2 - (9 x 109) q4/(0.25)2

1.8 x 106 = (9 x 109) (9 x 10-6)/(0.07)2 - (9 x 109) q4/(0.25)2

q4 = 102.3 x 10-6 C

when q4 is negative :

Ex = E3 - E4 = k q3/r32 + k q4/r42 = (9 x 109) (9 x 10-6)/(0.07)2 + (9 x 109) q4/(0.25)2

1.8 x 106 = (9 x 109) (9 x 10-6)/(0.07)2 + (9 x 109) q4/(0.25)2

q4 = - 102.3 x 10-6 C